OpenStudy (anonymous):

If f(x)= 2a|3x – 9| – ax, where a is some constant, find f ′(3).

OpenStudy (anonymous):

ok there are 2 cases 1 case is that the absolute is - and the othere is that its +

OpenStudy (anonymous):

a. 0 b. DNE c. 1 d. not enough information

OpenStudy (anonymous):

i have no idea wht u mean

OpenStudy (anonymous):

@iambatman @eliassaab @Michele_Laino

OpenStudy (anonymous):

they're asking for f' the derivative right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where we subs. 3 for x

OpenStudy (anonymous):

i just dont know how to do it because when we first find the derivative of x its a single x the x turns into a 1

OpenStudy (anonymous):

@ganeshie8 @eliassaab please help

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left| f \left( x \right) \right|=\frac{ f \left( x \right) }{ \left| f \left( x \right) \right| }f \prime \left( x \right)\]

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

\[f \prime \left( x \right)=2a \frac{ 3x-9 }{ \left| 3x-9 \right| }*3-a\]

OpenStudy (anonymous):

have you followed?

OpenStudy (anonymous):

doing it rn

OpenStudy (anonymous):

2a * 3-a

OpenStudy (anonymous):

???

OpenStudy (anonymous):

it is not derivable at x=3 because denominator is zero

OpenStudy (anonymous):

ohh so dne

OpenStudy (anonymous):

but wait

OpenStudy (anonymous):

wht about the a's doesnt it make this equation not solvable because we dont know wht the value of a is?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

no \[\left| x \right| ~is~continuous~at~x=0~but~\not~derivable~at ~x=0\] \[\left| x-a \right|~is~continuous~but~\not~derivable~at~x=a,whre~a~is~a~constant.\]

OpenStudy (anonymous):

so it dne?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

yes,correct because 3x-9=3*3-9=0

OpenStudy (anonymous):

but zero is also an answer

OpenStudy (anonymous):

a. 0 b. DNE c. 1 d. not enough information

OpenStudy (anonymous):

i have shown you \[\left| f \left( x \right) \right|=0~at~x=3\]

OpenStudy (anonymous):

so option b is correct as you have mentioned.

OpenStudy (anonymous):

thanks for the help :)

OpenStudy (anonymous):

yw

myininaya (myininaya):

if a=0 though then f(x)=0 we don't know what a is

myininaya (myininaya):

so when a is not equal to zero then f'(3) dne

myininaya (myininaya):

i'm saying I think we don't have enough information